daniela926:

a printer will produce 80 wedding invitations for $210. the price to produce 120 invitations is $290. the printer uses a linear function to dtermine the price for producing diff amounts of invitations. how much iwill the printer charge to produce 60 invitations

kittybasil:

Is there more to the question?

Jaynator495:

... ok? Good for the printer? Or is there a question to go along with that?

kittybasil:

If you're looking for "how many invitations will the printer produce in \(x\) amount of time", all you have to do is: (If this is "80 invitations per hour")\[\frac{80\text{invitations}}{1\text{hour}}\times\frac{x\text{hours}}{--}\]

kittybasil:

That was an example btw. Please update with the actual details or we can't help you ಠ_ಠ

daniela926:

thats the whole question

jhonyy9:

120 invit


$290 120-40=80 invit.
$210 = $290 -$80 80-40=40 invit.
$210-$80=$130 this is the price of 40 invitations 40-20=20 invit
$80-$40=$40 this is the price of 20 invit. the price of 60 invit. = $130 +$40 = $170